OpenStudy (dancergirl45):

Solve the system of equations. 5x + y = 9 3x + 2y = 4 A (−2, 5) B (1, 4) C (2, −1) D (4, −4)

pooja195 (pooja195):

Ok number the equation first 5x + y = 9 <---equation 1 3x + 2y = 4 <----equation 2 First we need to isolate a varible (so we can do substituion) 5x =5x+ y = 9 -5x When you do subtract 5x from both sides you get: y=-5x+9 Now we can substitute that into E2 \[\huge~\rm~\bf3x + 2(-5x+9)= 4\]Solve for x

OpenStudy (anonymous):

There are so many of these... I'm pretty sure that most people are too lazy to plug in the values lol

OpenStudy (dancergirl45):

-2X+11=4

pooja195 (pooja195):

So we are solving for x  Distribute 2 into the parehteses

pooja195 (pooja195):

What is 2 times -5?

OpenStudy (dancergirl45):

-10

pooja195 (pooja195):

Good what is 2 x 9?

OpenStudy (dancergirl45):

18

pooja195 (pooja195):

Good so now we have \[\huge~\rm~3x−10x+18=4\] Now we need to combine like terms \[\huge~\rm~\bf 3x-10x=?\]

OpenStudy (dancergirl45):

7X

pooja195 (pooja195):

* -7x 3-10=-7 Now this is what we have \[\huge~\rm~\bf−7x+18=4 \] We need to subtract 18 from both sides \[\huge~\rm~\bf−7x+18-18=4-18 \] \[\huge~\rm~\bf−7x=4-18 \] 4-18=?

OpenStudy (dancergirl45):

-14

pooja195 (pooja195):

Perfect! :) \[\huge~\rm~\bf−7x/-7=−14/-7\] Now all we have to do is divide both sides by -7 x=?

OpenStudy (dancergirl45):

-7X,-14

pooja195 (pooja195):

-7/-7=1 so just x What is -14/-7?

OpenStudy (dancergirl45):

2

OpenStudy (dancergirl45):

SO c

pooja195 (pooja195):

Yep =)

OpenStudy (dancergirl45):

tHANK YOU SO MUCH CAN U HELP ME WITH ONE MORE i WILL OPEN A NEW QUESTION

pooja195 (pooja195):

Okie ^-^ and yw